# Ting En Lam1,2, Yuhan Chen1,3, Elston Tan1,4, Eric Peh1, Ruirui Chen1, Paritosh Parmar1, Basura Fernando1

### v1

@@ Abstract
Causal video question answering (QA) has garnered increasing interest, yet existing datasets often lack depth in causal reasoning analysis. To address this gap, we capitalize on the unique properties of cartoons and construct CausalChaos!, a novel, challenging causal Why-QA dataset built upon the iconic “Tom and Jerry" cartoon series. With thoughtful questions and multi-level answers, our dataset contains much longer causal chains embedded in dynamic interactions and visuals, at the same time principles of animation allows animators to create well-defined, unambiguous causal relationships. These factors allow models to solve more challenging, yet well-defined causal relationships. We also introduce hard negative mining, including CausalConfusion version. While models perform well, there is much room for improvement, especially, on open-ended answers. We identify more advanced/explicit causal relationship modeling & joint modeling of vision and language as the immediate areas for future efforts to focus upon. Along with the other complementary datasets, our new challenging dataset will pave the way for these developments in the field. We will release our dataset, codes, and models to help future efforts in this domain.
Figure 1: (Left) Examples of causal questions about characters’ actions from our CausalChaos! dataset—based on Tom & Jerry cartoon series. Q: question; A: answer; E: explanation. Please view in Adobe Reader to play the embedded videos for better explanation. (Right, top) Illustration of causal chain, scene changes. Linking multiple clues/cues embedded in different scenes to solve causal relationships pose a challenge for VideoQA models. (Right, bottom) Being a synthetic medium allows animators to leverage Principles of Animation to stylize the visuals & motions to disentangle/highlight key content of the scene to create well-defined/unambiguous and effectively communicated cause-and-effect relationships. The interplay of these factors allow models to focus on solving complex, yet, well-defined, unambiguous causal relationships.
@@ 1 Introduction
Causal video question answering has been gaining increasing interest [52, 28, 47, 17]. However, our analysis, as presented in Figs.1, 2, and in § 2, indicates that existing causal QA datasets may often require surface-level understanding or involve more of a simple word substitution in the QA pairs, rather than causal reasoning (e.g., “Q. Why are the cars on the street not moving? A. Parked."). Consequently, these datasets do not demand true causal reasoning analysis. This gap leaves a dearth of reliable benchmarks for evaluating the causal reasoning capabilities of video understanding models. Hence, in this study, we concentrate on developing a rigorous and challenging causal Video-QA dataset. We aim to create a high-quality resource that can effectively test and facilitate the advancement of sophisticated causal models.
@@
  • As illustrated in Figs.1, 2, our emphasis lies in crafting thoughtful questions paired with multi-layered answers, aiming to cultivate a deeper and more comprehensive grasp of causal relationships. Compared to existing datasets, CausalChaos! presents extended chains of causality [^1] A causal chain is a sequence of events/actions in which each step influences or leads to the next, creating a cause-and-effect relationship. It illustrates how one event/action produces a subsequent event/outcome., as highlighted in Fig.2 and discussed in § 2.
  • Video clips in our dataset feature frequent scene and shot changes [^2] A scene change or shot change refers to the transition from one visual setting or perspective to another. It typically involves shifting focus to a new location, action, or character within the storyline, often signaled by techniques such as cuts & fades.. The links or intermediary causes within the causal chains are often dispersed across various scenes. Consequently, models are challenged to exert greater cognitive effort in connecting multiple events (scenes) and identifying intermediate causes to comprehend the “why" questions related to the clips.
  • Despite the complexity and length of the causal chains, they are distinctly delineated, devoid of ambiguity, and communicated effectively using principles of animation [42] such as staging and exaggeration. This deliberate design allows models to concentrate on deciphering causal relationships.
@@
  • Our dataset demands a diverse range of reasoning skills, encompassing deductive, spatial, emotional reasoning, and more, as outlined in Fig.2 and discussed in §3.
We evaluated various state-of-the-art Video-QA models including recent multimodal instruction tuning models and found that our dataset remains the most challenging causal QA dataset. Evaluating on our dataset, helped us reveal that models generally jumped to conclusions based on partial evidence, instead of taking into consideration full set of evidence; models failed to engage in true causal reasoning, trying to select answers based on shortcuts like object/action-noun/verb matching. Based on this, we identified more advanced/explicit causal relationship modeling and joint modeling vision and language as the immediate areas for future efforts to focus upon. Further, we show that similar to how cartoons help children better connect cause and effects, they can help Video-QA models as well. We incorporated our dataset with the NextQA [47], a real-world dataset and found that it brings some improvements in why questions. What is more, incorporating our dataset brings improvement on non-Why questions as well.
Overall, we believe that our CausalChaos! dataset contributes challenges that span the entire Video-QA pipeline, from deciphering intricate videos to processing complex questions and discerning between nuanced answers, stimulating research in many areas like video processing, causal reasoning, language modeling and joint modeling. Along with the other complementary datasets, such as [47, 17, 19, 21, 46], our new challenging dataset will pave the way for these developments in the field. We advocate for leveraging our dataset sourced from cartoons to inform the model design process. Our dataset better reflects the challenges Video-QA models may face in the real world, such as longer causal chains and frequent scene transitions in visual streams. Nevertheless, we caution against blindly transferring or deploying entire models/weights trained on our dataset to real-world applications.
@@
Multilevel
Answers
Causal
Chain Len
Scene
Reasoning
Spectrum
Multigranular
Actions
NextQA [47] 1 1 narrow limited
CausalVidQA [17] 1 1 narrow limited
IntentQA [19] 1 1 narrow limited
Ours 2.7 4 wide ubiquitous
@@
Figure 2: (a) Types of reasoning demanded by our CausalChaos! dataset. Reasoning types: DR-deductive reasoning; IR-inductive; SR-spatial; CR-causal; CT-critical thinking; ER-emotional; AR-abductive; TR-temporal; None-no reasoning required as per the human subjects. None is undesirable, and tend to indicate that questions are less challenging. (b) Comparison among CausalChaos! and existing causal videoQA datasets. (c) Qualitative comparison between CausalChaos! and NextQA datasets. CausalChaos! Answers and Explanations give detailed information regarding the actual cause-and-effect relationships, motivations, emotions covering wide range of reasoning types. Note that, here, with the help of human annotators, we have temporally cropped videos to retain only the relevant parts from NextQA dataset videos; otherwise, raw videos are longer resulting in unintended problem of temporal localization for models.
@@ 2 Related Work
VideoQA Datasets and Benchmarks. To drive the progress in VideoQA, researchers have developed various datasets with distinct focuses and contributions—see the survey [59]. In the following, we mainly compare and contrast our work from prior art, with a focus on causal QA literature. We provide a detailed discussion on Video QA datasets in the Supplementary material.
Early VideoQA datasets
such as MovieFIB [26], MSVD-QA [48], MSRVTT-QA [48], YouTube2Text [10], open-ended QA [58], Zeng et al. [56], and Video Context-QA [61] relied on video captions or descriptions to automatically generate questions and corresponding answers, primarily focusing on object and action recognition. However, they lacked depth in understanding complex interactions and causal relationships within videos. TGIF-QA [13, 14] and ActivityNet-QA [54], while covering short and long web videos respectively, still fell short in capturing comprehensive reasoning, particularly in object interactions and causal relationships. Social-IQ [55] addressed human social behaviour in videos through multimodal cues but lacked broader context and visual reasoning capabilities. AGQA [9] focused on spatio-temporal scene understanding, but did not cover explanatory "why" questions. STAR [46], a situation reasoning dataset, extended to prediction and feasibility questions but also lacked explanatory inquiries.
Causal QA datasets.
Our work is inspired from a few datasets: CLEVRER [52, 28], NextQA [47], CausalVidQA [17], IntentQA [19] that cover causal reasoning. However, they have several limitations that our dataset aims to fill:
@@
  • Some are limited by their scope, e.g., [52, 28] only consider collision events of simple inanimate objects such as cuboids; as a result, lack: actors, their characteristics such as emotions, intentions, lack diverse language descriptions.
@@
  • Some datasets (e.g., [47, 17]) do not contain any precise temporal annotations, which inadvertently involves temporal localization & may create a false sense of difficulty, while the questions may not be challenging. We alleviate this problem by providing precise temporal annotations and focusing on designing truely complex and challenging QA.
@@
  • Some datasets lack complexity in questions [47, 17, 19]. Based on human studies, we found that a majority of causal QA in these datasets were flagged as not requiring any notable reasoning (Figure 2). We also attempted to quantify the complexity in reasoning by computing the lengths of causal chains involved in their QAs vs. ours using GPT-4 [32]. Causal chains in [47, 17, 19] have an average length of 1. In comparison, our dataset has much longer causal chains—average length 2.7—posing a challenge for models in connecting multiple events or cues. Our dataset also demands wide spectrum of types of reasoning (2(a)).
@@
  • Scenes in [52, 28, 47, 17, 19] are less dynamic. For example, they mostly involve a single scene. In comparison, our dataset, on average, has about 4 scene changes. Rapid scene changes & dynamic interactions pose a challenge to VideoQA models when modeling cause-&-effect relationships over longer causal chains by linking context and cues embedded in different scenes (more details in subsection 3.2). While related challenges have been acknowledged in other computer vision problems [34], they remain unexplored in the realm of VideoQA.
@@
  • Current datasets do not emphasize on hard incorrect option selection. As such, models may not require to do causal reasoning, but rather they can exploit shortcuts like object/action-noun/verb matching in vision-language spaces to correctly answer. Thus, we develop hard negative selection strategies.
@@
  • These datasets contain a single-level answer to ‘Why’-questions. On the other hand, we richly annotate our dataset to provide two-level answers to ‘Why’-questions. Quality of answers is further enhanced, and are more informative as a consequence of the more complex questions.
VideoQA Tasks/Models.
VideoQA generally uses two protocols: 1) multichoice QA (MCQA) & 2) open-ended answer generation (OEAG). MCQA involves choosing the correct option from the multiple provided options. MCQA models [2, 8, 6, 16, 15, 7] fuse Video, question, & answer representations & process them using a multi-layer perceptron, which ultimately predicts the correct answer. On the other hand, in much more challenging OEAG, VideoQA models generate answers in natural language [2, 51, 6, 15]. This is typically accomplished by decoding the fused video & question representations into a sequence of words (answer). Recently, vision-language and multimodal instruction tuning models [7, 18, 57, 21] have shown strong performance on video understanding tasks. We evaluate & report the performance of all these models on our dataset. Further discussion on individual models can be found in Supplementary material.
@@ 3 CausalChaos! Dataset
In this section, we detail the construction of our challenging VideoQA dataset focused on causal reasoning. We provide insights into the video source, annotation process, and rigorous quality checks employed to ensure detailed annotations of high quality. Further, we discuss the unique attributes of our dataset that make it a valuable resource for the causal action/VideoQA task.
@@ 3.1 Dataset Construction
To exclusively focus on visual information for reasoning, we selected the timeless silent cartoon series, ‘Tom & Jerry’ (1940), spanning over 6 seasons and encompassing a total of 161 episodes. Silent cartoons that do not contain modalities other than vision (such as dialogues) can foster the development of visual reasoning capabilities in children. Similarly, we hypothesize that using silent cartoons for VideoQA can allow for a focus on visual reasoning. The Tom & Jerry series is also an ideal source for constructing a causal reasoning dataset, as it provides abundant segments showcasing diverse cause-&-effect relationships. A concurrent work [22] on causal image generation has also leveraged Tom & Jerry cartoon.
@@
We collect the following annotations.
$\bullet$ Questions. The questions are crafted to capture the why or reasoning behind the actions of characters in Tom & Jerry cartoons. To cultivate comprehensive and deeper video understanding capabilities in models, we formulated thoughtful causal questions where cause & effect are connected by longer causal chains. Questions also extend beyond explicit visual cues, encompassing gestures & expressions, to delve into the characters’ underlying intentions and goals.
The annotators would generate questions while watching the video for the first time to ensure that they did not have prior knowledge of the scene, which mimics how models would assess unseen clips. Subsequently, they would rewatch the videos multiple times to generate more critical thinking questions.
While generating questions, annotators watch the videos without audio to ensure that the dataset does not contain audio cues to focus on visual reasoning.
$\bullet$ Multi-level Answers. The first-level or the Primary answer, represents a shallower, literal or direct cause or form of response. It is accompanied by a deeper form of Explanation, which considers the broader context of the scene, the thoughts, feelings, intentions of characters, & their actions. This deeper explanation also takes into account potential consequences & provides further reasoning to support the primary answer. It includes reasons & additional details to comprehensively address the question. Examples shown in Figure 1 & Supplementary material.
$\bullet$ Temporal annotations. For each QAE set the Start &End frame numbers of the video are recorded such that the entire scene including frames that offer context & details supporting the reasons behind the actions is covered.
In summary, each dataset sample is annotated with: {start frame, end frame, Question, Answer, Explanation}.
@@
To ensure the quality and reliability of the dataset, we implemented a two-stage quality check process. In the first stage, quality checkers conducted thorough assessments of episodes they did not annotate. Subsequently, in the second stage, quality checkers reviewed episodes for which they did not perform the first-stage quality check.
Multiple quality checkers were involved in the process, carefully reviewing and verifying curated question and answer sets for logical fallacies, timestamp inconsistencies, grammatical errors, and spelling mistakes. The process also involved checking and excluding annotations containing audio or text aspects that the model cannot comprehend. To prevent direct overwriting, any edits were flagged in a different font color for further discussion.
The inconsistencies or errors that were flagged out were resolved through discussions and consensus among the checkers and respective annotators. In these discussions, the team referred to the episode synopses available on Tom and Jerry Wiki [43], an online Tom and Jerry encyclopedia. This external source served as a third-party opinion and facilitated fact-checking. By incorporating multiple perspectives, the discussions helped to provide a deeper level of analysis to generate a more challenging and comprehensive dataset.
@@
Annotators meticulously watched all 161 cartoon episodes spanning 6 seasons, generating a total of 4945 richly annotated Question-Answer-Explanation sets for the dataset. The inclusion of a substantial number of Question-Answer-Explanation sets across all seasons, covering a wide range of scenes, characters, and events, underscores the comprehensive nature of our dataset. This diversity establishes a robust foundation for understanding and answering questions related to the cartoon series. Word clouds corresponding to answers and explanations from our dataset are presented in Supplementary material.
3.2 Unique Characteristics of our dataset

### v2

# Yuhan Chen1,3, Elston Tan1,4, Basura Fernando1 1

@@ Abstract
Causal video question answering (QA) has garnered increasing interest, yet existing datasets often lack depth in causal reasoning. To address this gap, we capitalize on the unique properties of cartoons and construct CausalChaos!, a novel, challenging causal Why-QA dataset built upon the iconic “Tom and Jerry" cartoon series. Cartoons use the principles of animation that allow animators to create expressive, unambiguous causal relationships between events to form a coherent storyline. Utilizing these properties, along with thought-provoking questions and multi-level answers (answer and detailed causal explanation), our questions involve causal chains that interconnect multiple dynamic interactions between characters and visual scenes. These factors demand models to solve more challenging, yet well-defined causal relationships. We also introduce hard incorrect answer mining, including a causally confusing version that is even more challenging. While models perform well, there is much room for improvement, especially, on open-ended answers. We identify more advanced/explicit causal relationship modeling & joint modeling of vision and language as the immediate areas for future efforts to focus upon. Along with the other complementary datasets, our new challenging dataset will pave the way for these developments in the field.
@@ 1 Introduction
Understanding the intricate motivations behind human actions is paramount in developing sophisticated systems capable of nuanced behavior analysis. In real-world scenarios, actions are shaped by a multitude of factors, including personal experiences, emotions, social contexts, and cultural backgrounds. This complexity necessitates a comprehensive approach to unraveling the "why" behind actions, fostering empathy, effective communication, and robust decision-making. Causal video question answering aims to decipher the answers behind characters’ actions. Despite the growing interest in causal video QA, existing datasets often fall short, requiring only: 1) surface-level understanding; or 2) involve more of simple word substitution in the QA pairs, rather than causal reasoning (e.g., “Q. Why are the cars on the street not moving? A. Parked."). Recognizing this gap, we embark on the development of a rigorous and challenging causal Video-QA dataset. Our goal is to provide a high-quality resource that rigorously evaluates and advances causal video QA models.
Figure 1: (Left) Examples of causal questions about characters’ actions from our CausalChaos! dataset—based on Tom & Jerry cartoon series. Q: question; A: answer; E: explanation. Please view in Adobe Reader to play the embedded videos for better explanation. (Middle) Illustration of causal chain, scene changes. Linking multiple clues/cues embedded in different scenes to solve causal relationships pose a challenge for VideoQA models. (Right) Animators leverage Principles of Animation to stylize the visuals & motions to disentangle/highlight key content of the scene to create well-defined/unambiguous and effectively communicated cause-and-effect relationships. The interplay of these factors allow models to focus on solving complex, yet, well-defined, unambiguous causal relationships.
@@
  • As demonstrated in Figure 1, 3, we formulate thought-provoking questions, where the answer (A) and its detailed explanation (E) aim to enhance the model’s understanding of the causal chains of visual events in a given video clip. A causal chain is a sequence of events/actions in which each step influences or leads to the next, creating a cause-and-effect relationship. It illustrates how one event/action produces a subsequent event/outcome. Compared to existing datasets, CausalChaos! presents longer causal chains, as highlighted in Figure 3 and discussed in section 2.
  • Video clips in our dataset feature frequent scene and shot changes. Here, a scene change or shot change refers to the transition from one visual setting or perspective to another. It typically involves shifting focus to a new location, action, or character within the storyline. The links and causes of the causal chains are often dispersed across various scenes. Consequently, models are challenged to exert greater cognitive effort in connecting multiple events (scenes) and identifying intermediate causes to comprehend the “why" questions related to the clips.
  • Despite the complexity and length of the causal chains, they are distinctly delineated, unambiguous, and effectively communicated using principles of animation [42] like staging and exaggeration. This deliberate design allows models to focus on deciphering causal relationships.
@@
  • Our dataset demands a diverse range of reasoning skills, encompassing deductive, spatial, emotional reasoning, and more, as outlined in Figure 3 and discussed in section 3.
Upon evaluating various state-of-the-art VideoQA models including the recent multimodal instruction tuning models, we found that our dataset remains one of the most challenging causal QA dataset. Particularly, we observed that models often: 1) jumped to conclusions based on partial evidence, rather than considering the full set of evidence; 2) failed to engage in true causal reasoning, opting instead for shortcuts like object/action-noun/verb matching to select answers. Based on this, we identified more advanced/explicit causal relationship modeling and jointly modeling vision and language as the immediate areas for future efforts to focus upon. Further, we show that similar to how cartoons help children better connect cause and effects, they can help VideoQA models as well. We incorporated our dataset with the NextQA [47], a real-world dataset and found that it brings some improvements in why questions. What is more, incorporating our dataset brings improvement on non-Why questions as well.
Overall, we believe our CausalChaos! dataset presents challenges spanning the entire VideoQA pipeline, from deciphering intricate videos to processing complex questions and discerning nuanced answers, stimulating research in video processing, causal reasoning, language modeling, and joint modeling. Along with the other complementary datasets, such as [47, 17, 19, 21, 46], our new challenging dataset will pave the way for these developments in the field.
@@
MA
CCL
NOS
RS
MGA
NextQA 1.1 1 narrow limited
CausalVid 0.9 1 narrow limited
IntentQA 1.3 1 narrow limited
CausalChaos! 1.9 4 wide abundant
@@
Figure 3: (a) Types of reasoning demanded by our CausalChaos! dataset. Reasoning types: DR-deductive reasoning; IR-inductive; SR-spatial; CR-causal; CT-critical thinking; ER-emotional; AR-abductive; TR-temporal; None-no reasoning required as per the human subjects. None is undesirable, and tend to indicate that questions are less challenging. (b) Comparison among CausalChaos! and existing causal videoQA datasets. MA-multilevel answers; CCL-causal chain length; NOS-no. of scenes; RS-reasoning spectrum; MGA-multigranular actions. (c) Qualitative comparison between CausalChaos! and NextQA (Why-QA) datasets. CausalChaos! Answers and Explanations give detailed information regarding the actual cause-and-effect relationships, motivations, emotions covering wide range of reasoning types. Note that, we have temporally cropped videos to retain only the relevant parts from NextQA dataset videos; otherwise, raw videos are longer resulting in unintended problem of temporal localization for models.
@@ 2 Related Work
To drive the progress in VideoQA, researchers have developed various datasets [26, 48, 10, 58, 56, 61, 13, 14, 54, 55, 9, 46] with distinct focuses and contributions—see the survey [59]. In the following, we mainly compare our work with prior art focusing on causal QA literature. We provide a detailed discussion on various VideoQA datasets and models in the Appendix. Our work is inspired from a few datasets: CLEVRER [52, 27], NextQA [47], CausalVidQA [17], IntentQA [19] that cover causal reasoning. However, they have limitations that our dataset aims to fill:
@@
  • Limitation in scope: e.g., [52, 27] only consider collision events of simple inanimate objects such as cuboids; as a result, lack: actors, their characteristics such as emotions, intentions. Our dataset has actors whose actions are shaped by emotions, intentions, context, etc.
@@
  • Lack of precise temporal annotations (e.g., [47, 17]) inadvertently involves temporal localization & may create a false sense of difficulty, while the questions may not be challenging. We alleviate this problem by providing precise temporal annotations and focusing on designing truly complex and challenging QA.
@@
  • Lack of complexity in questions (e.g.[47, 17, 19]). Based on human studies, we found that a majority of causal QA in these datasets were flagged as not requiring any notable reasoning (Figure 3(a)). We also attempted to quantify the complexity in reasoning by computing the lengths of causal chains involved in their QAs vs. ours using GPT-4o [32]. Causal chains in [47, 17, 19] are shorter (Figure 3(b). In comparison, our dataset has longer causal chains—average length 1.9—posing a challenge for models in connecting multiple events or cues. Our dataset also demands wide spectrum of types of reasoning (3(a)).
@@
  • Scenes in [52, 27, 47, 17, 19] are less dynamic, mostly involving a single scene. In comparison, our dataset averages about four scenes. Rapid scene changes & dynamic interactions challenge VideoQA models to link context & cues across different scenes, modeling cause-and-effect relationships over longer causal chains (more details in subsection 3.2). While similar challenges have been acknowledged in other computer vision problems [34], they remain unexplored in VideoQA.
@@
  • Lack of hard negatives. Current datasets may not emphasize on including hard incorrect options. Due to this, models may not be required to do causal reasoning, but rather they can exploit shortcuts like object/action-noun/verb matching in vision-language spaces to select correct answer. To address this limitation, we develop hard negative selection strategies.
@@
  • No hierarchical answers with different level of explanation. These datasets contain a single-level answer to ‘Why’-questions. On the other hand, we richly annotate our dataset to provide two-level answers to ‘Why’-questions—1) direct/immediate cause; 2) deeper explanation. Quality of answers is further enhanced, and are more informative as a consequence of the more complex questions.
@@ 3 CausalChaos! Dataset
This section details the construction of our VideoQA dataset designed to challenge causal reasoning, covering its video source, annotation process, & quality checks to ensure high-quality annotations. We then discuss the unique attributes that make our dataset valuable for causal VideoQA tasks.
@@ 3.1 Dataset Construction
To focus on visual reasoning, we selected the timeless silent cartoon series, “Tom & Jerry" (1940) [^1] We only provide the annotations; Videos can be obtained from: https://www.warnerbros.com/tv/tom-and-jerry-1965-volume-1., spanning over six seasons and 161 episodes. Silent cartoons, which lack other modalities like dialogue, foster visual reasoning in children. Similarly, we hypothesize that using silent cartoons for VideoQA will enhance visual reasoning. The Tom & Jerry series is ideal for a causal reasoning dataset, offering abundant segments with diverse cause-and-effect relationships. Concurrent work [22] on causal image generation also leverages Tom & Jerry.
@@
Each dataset sample includes the following annotations: {start frame, end frame, Question, Answer, Explanation}. Details in the following.
$\bullet$ Questions are crafted to capture the why or reasoning behind the actions of characters in Tom & Jerry cartoons. To cultivate comprehensive & deeper video understanding capabilities in models, we formulated thought-provoking causal questions where cause & effect are connected by longer causal chains. Questions also extend beyond explicit visual cues, encompassing gestures & expressions, to delve into the characters’ underlying intentions & goals. Annotators generated questions while watching the video for the first time to mimic how models assess unseen clips. They then rewatched the videos multiple times to create more critical thinking questions. To focus on visual reasoning, annotators watched the videos without audio, ensuring no audio cues were included in the dataset.
$\bullet$ Multi-level Answers. The first-level or the Primary answer, represents a literal or direct cause or form of response. It is accompanied by a deeper form of Explanation, which considers the broader context of the scene, the thoughts, feelings, intentions of characters, & their actions. This deeper explanation also takes into account potential consequences & provides further reasoning to support the primary answer. It includes reasons & additional details to comprehensively address the question. Examples shown in Figure 1 & Appendix.
$\bullet$ Temporal annotations. For each QAE set, the Start and End frame numbers are recorded to cover the entire scene, including contextual frames that support the reasons behind the actions.
@@
To ensure dataset quality and reliability, we implemented a two-stage quality check process. First, quality checkers assessed episodes they did not annotate. Then, they reviewed episodes they did not check in the first stage. Multiple quality checkers carefully reviewed and verified curated question and answer sets for logical fallacies, timestamp inconsistencies, grammatical errors, and spelling mistakes. They also excluded annotations containing audio or text aspects that the model cannot comprehend. To prevent direct overwriting, edits were flagged in a different font color for further discussion. Flagged inconsistencies or errors were resolved through discussions and consensus among checkers and annotators. The team referred to episode synopses from Tom and Jerry Wiki [43], an online encyclopedia, as a third-party opinion and for fact-checking.
@@
Annotators reviewed all 161 episodes across 6 seasons, generating 4,945 detailed Question-Answer-Explanation sets. This comprehensive dataset covers various scenes, characters, and events, providing a solid foundation for understanding and answering questions about the series. Additional statistics and word clouds for answers and explanations are available in the Appendix.
Figure 4: Grounded in diverse visual information.
Figure 5: Traditional temporal modeling vs Dynamic scene linking. Notice the abrupt scene change, which causes disruption in visual flow, resulting in large amplitude and widespread optical flow.
3.2 Unique Characteristics of our CausalChaos! Dataset